OpenStudy (leahhhmorgannn):

The function H(t) = -16t2 + vt + s shows the height H (t), in feet, of a projectile launched vertically from s feet above the ground after t seconds. The initial speed of the projectile is v feet per second. Part A: The projectile was launched from a height of 82 feet with an initial velocity of 60 feet per second. Create an equation to find the time taken by the projectile to fall on the ground. (2 points) Part B: What is the maximum height that the projectile will reach? Show your work.

OpenStudy (leahhhmorgannn):

Part C: Another object moves in the air along the path of g(t) = 10 + 63.8t where g(t) is the height, in feet, of the object from the ground at time t seconds. Use a table to find the approximate solution to the equation H(t) = g(t), and explain what the solution represents in the context of the problem? [Use the function H(t) obtained in Part A, and estimate using integer values] (4 points) Part D: Do H(t) and g(t) intersect when the projectile is going up or down, and how do you know? (2 points)

OpenStudy (leahhhmorgannn):

I've seen people say they need A in order to help, which I've already found A. Part A: -16t^2 + 60t + 82

OpenStudy (love_ranaa):

Sorry, but I have no idea ..

OpenStudy (dmndlife24):

I guess this is how you would do it... A H(t)=-16t^2+60t+82 You replace H(t) with 0 to find when it falls on the ground. B H'(t)=v(t)=-32t+60 t=15/8 H(15/8)=138.25 ff